Find the derivative of the function. 4x - 7 8x + 1 y
Answer #1

@ 163-8)9-da- substitute = 6x²5 de esta lor 2-s] = 6(2x) -0 = 128. de = it a I 23 Ya hayo? Le stall dren] gew=fres (go? (gen)